OpenStudy (theghostlycreeper):

Turn 12% to a fraction and a decimal.

OpenStudy (tootsi123):

12/100 would be your fraction

OpenStudy (tootsi123):

the decimal would be 0.12...all you do is move the decimal over two to the left

OpenStudy (onyxxelite):

12/100, .12

OpenStudy (mysticfury):

^^^^ they are correct, you can just simply put the percent over 100 to get your fraction and for a decimal all you have to do is move it over 2 places to the left :)

OpenStudy (theghostlycreeper):

can you help me with another @tootsi123

OpenStudy (mathstudent55):

The symbol % is called "percent". The word "percent" literally means "per 100", or "divided by 100". That means any percent can be changed to a fraction by simply setting it as a fraction over 100. Any percent can changed into a decimal by simply dividing it by 100. Dividing by 100 is just moving the decimal point two places to the left. Examples: Percent Fraction Decimal 10% = 10/100 = 0.10 25% = 25/100 = 0.25 78% = 78/100 = 0.78 300% = 300/100 = 3